Q Search this course Interest rate premiums Excel Online Structured Activity: Interest rate premiums A 5-year Treasury bond has a 3% yield. A 10-year Treasury bond yields 6.1%, and a 10-year corporate bond yelds 8 65%. average 3.6% over the next 10 years (IP10-3.6%). Assume that there is no maturity risk premium (MRP-0) The mar ket expects that inflation will er the next 10 years. (Hint: Remember that the default risk premium and the liquidity premium are zero for Treasury secuitis:DRPIP-0) A 5-year corporate bond has the same default t risk premium and liquidity premium as the 10-year corporate bond described. The data has been collected in the and Microsoft Excel Online file below. Open the spreadsheet and perform the required analysis to answer the question below Open spreadsheet What is the yield on this 5-year corporate bond? Round your answer to two decimal places Type here to search
Tell me what you want to File Home. Insert Data Review View Ξ Ξ percentage Arial L' Paste Numi Alignment Undo Clipboard Font 814 1 Interest rate premiums 3 15-year Treasury yeld (Ts) 4 10-year Treasury yield (T10) 5 10-year Corporate yield (C1o 6 300% 6.10% 8 65% |Inflaton Premium over 10 years (IP10) 7 Maturity Risk Premium (MRP) 8 DRP Treasury 9 LP Treasury 3 6096 000% 0 00% 00096 Formulas 12 Real risk-free rate,r 13 14 Inflation premium over 5 years (Ps) 15 17 18 5-year Corporate yield (Cs) 20 21 Sheet1 ü O Type here to search
Answer #1

Answer:

Risk Free Rate Rf= T10-IP10=6.10%-3.60%=2.50%

Inflation premium for 5 years IP5=T5-Rf=3.00%-2.50%=0.5%

DRP10+LP10=C10-T10=8.65%-6.10%=2.55%

Since DRP5+LP5=DRP10+LP10=2.55%

5 years corporate bond = Rf+IP5+DRP5+LP5=2.5%+0.5%+2.55%=5.55%